• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name suleiman

    Log entry time 08:00:08 on November30,2006

    Entry number 189560

    keyword=Owl Shift Summary


    - 00:00 - 01:00 No beam.
    - 01:00 - 01:45 beam is ready but cannot take it because target
    is unstable.
    - 01:45 5 uA on H2. Re-start Coda.
    - 01:56 Run 1146 Pion Rejector Calibration
    - 02:30 Paged Mike Seely - He is coming to put in a new H2 bottle.
    The target temperature is now 19.30 K and the pressure
    is 16.5 psia.
    - 02:45 Run 1147, 1148 Prl and Cerenkov Cal (changed Cerrenkov HV
    compared to the previous run).
    - 03:21 Run 1149 Prl and Cerenkov Q3 Left has positive polarity

    - 3:25-4:30 Controlled access to fill H2.

    - 4:30 Run 1150 back to ep elastic + Left at +4%.
    - 4:40 Target pressure is 22.4 psia and temperature is 19.00 K
    beam current is 20 uA.
    - Runs 1151 and 1152 with two different Left S1 L-PMT 1 HV.

    - 05:45 Changed Left to -4%. Re-start Coda (TS11 crashed).
    - 06:05 Run 1153, 1154 (-4% Left).
    - Run 1155 Test of epics variables
    - 6:36 Changed back to nominal. Cycled Q2/3 Quads on Left.
    - Run 1156 ep elastic (no beam most of the run).
    - Run 1157 ep elastic (no beam most of the run).